OpenStudy (btaylor):

Need help on how to solve inequalities. I know how to do them graphing, but how do I do it mathematically? 1. \[x^2 + 2x - 3 \le 0\] 2. \[\frac{2x-1}{3x-2}\le1\] 3. \[x^2 + x + 1 > 0\]

OpenStudy (btaylor):

@ajprincess @PaxPolaris @KingGeorge @dpaInc can you help me?

OpenStudy (anonymous):

I do it by completing the square first. That is: x^2 + 2x -3 <= 0 <=> (x^2 + 2x + 1) - 3 - 1 <= 0 <=> (x+1)^2 <= 4 Now take the square root to get: |x+1| <= 2 There are 2 cases: i) If x+1 >= 0, then the modulus is: |x+1| = x+1 and we have x+1<=2 or x<=1. ii) if x+1 < 0, then |x+1| = -(x+1) and we have -(x+1) <=2 or x>= -3. So the solution will be: -3 <= x <= 1.

OpenStudy (anonymous):

No. 3 is easier to solve: x^2 + 2(1/2)x + (1/2)^2 - (1/2)^2 + 1 > 0 <=> (x+1/2)^2 > -3/4 The square on the left side is always positive so this inequality holds for all x.

OpenStudy (paxpolaris):

1. \[\Large x^2+2x-3 \le 0\]\[\implies \Large \left( x+3 \right)\left( x-1 \right) \le 0\] you can see that the graph of y=(x+3)(x-1) will look like this: so, from the graph \[\Large -3 \le x \le 1\]

OpenStudy (paxpolaris):

even without graphing this: for the product (x+3)(x-1) to be negative one of them has to be negative and the other positive so, \(\large x +3 \ge 0\), and \(\large x -1 \le 0\)

OpenStudy (paxpolaris):

and 2. \[\large {2x-1 \over 3x-2} \le 1\] \[\implies \Large 2x-1 \le 3x-2\ \dots\ \left\{ \text{ where } 3x-2 \ne0\right\}\]